OpenStudy (anonymous):

if f(x) = x^3+Ax^2+Bx-3 and if f(1) = 4 and f(-1) = -6 what is the value of 2A+B?

OpenStudy (anonymous):

f(1) = 1 + A + B - 3 = A + B -2 = 4 f(-1)= -1 + A -B -3 = A- B -4 = -6 A+B = 6 A-B = -2 2A = -4 A = -2 for A=-2, B = 8 2A + B = 2(-2) + 8 = 4

OpenStudy (anonymous):

\[f(1)=1+A+B-3=4\] so \[A+B+6\] \[f(-1)=-1+A-B-3=-6\] so \[A-B=-2\]

OpenStudy (anonymous):

solve \[A+B=6\] \[A-B=-2\] get \[A=2\] \[B=4\]

OpenStudy (anonymous):

oh right, what jyouie said!

OpenStudy (anonymous):

yours made more sense to me :P

OpenStudy (anonymous):

oh I've did something wrong...sorry...I mistakenly saw it as -6 +2 sorry though